# Is Chain-of-Thought Reasoning of LLMs a Mirage? A Data Distribution Lens

### v3

@@ A.1 Proof of CoT Generalization Bound
Let $f_{\theta}$ be a model trained on samples from the distribution $\mathcal{D}_{\text{train}}$ using a loss function $\ell(f_{\theta}(x),y)$ that is $\Lambda$-Lipschitz and bounded. The expected test risk is given by
$R_{\text{test}}(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{test}}}\left[\ell(f_{\theta}(x),y)\right].$ (15)
We can decompose the test risk as
$R_{\text{test}}(f_{\theta})=R_{\text{train}}(f_{\theta})+\left(R_{\text{test}}(f_{\theta})-R_{\text{train}}(f_{\theta})\right).$ (16)
To bound the discrepancy between $R_{\text{test}}$ and $R_{\text{train}}$, we invoke a standard result from statistical learning theory. Given that $\ell$ is $\Lambda$-Lipschitz and the discrepancy measure $\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$ is an integral probability metric (e.g., Wasserstein-1 distance), we have
$\left|R_{\text{test}}(f_{\theta})-R_{\text{train}}(f_{\theta})\right|\leq\Lambda\cdot\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}).$ (17)
Therefore, the test risk satisfies
$R_{\text{test}}(f_{\theta})\leq R_{\text{train}}(f_{\theta})+\Lambda\cdot\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}).$ (18)
We next account for the generalization gap between the empirical training risk $\hat{R}_{\text{train}}(f_{\theta})$ and the expected training risk $R_{\text{train}}(f_{\theta})$. By applying a concentration inequality (e.g., Hoeffding’s inequality), with probability at least $1-\delta$, we have
$R_{\text{train}}(f_{\theta})\leq\hat{R}_{\text{train}}(f_{\theta})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right),$ (19)
where $n$ is the number of training samples.
Combining the above, we obtain that with high probability,
$R_{\text{test}}(f_{\theta})\leq\hat{R}_{\text{train}}(f_{\theta})+\Lambda\cdot\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right).$ (20)
This concludes the proof. ∎
A.2 Proof of Task Generalization Failure Threshold
@@
$P(\text{correct}|C)=P_{0}\prod_{i=1}^{m}\rho_{a}^{\mathbb{I}[A_{i}]}\prod_{j=1}^{n}\rho_{f}^{\mathbb{I}[F_{j}]}\rho_{p}^{\mathbb{I}[\mathcal{Q}]}\rho_{c}^{C_{T}}$
@@
$\ln P(\text{correct}\mid C)\;=$ $\ln P_{0}+\sum_{i=1}^{m}\mathbb{I}[A_{i}]\,\ln\rho_{a}+\sum_{j=1}^{n}\mathbb{I}[F_{j}]\,\ln\rho_{f}+\mathbb{I}[\mathcal{Q}]\,\ln\rho_{p}+C_{T}\,\ln\rho_{c}$ (21)
@@
$\ln P(\text{correct}|C)=\ln P_{0}-\xi_{a}\sum_{i=1}^{m}\mathbb{I}[A_{i}]-\xi_{f}\sum_{j=1}^{n}\mathbb{I}[F_{j}]-\xi_{p}\mathbb{I}[\mathcal{Q}]-\xi_{c}C_{T}$ (22)
@@
$\ln P(\text{correct}|C)\leq\ln P_{0}-\delta\cdot\text{TGC}(C)$ (23)
@@
$\text{TGC}(C)=\alpha\sum_{i=1}^{m}\mathbb{I}[A_{i}]+\beta\sum_{j=1}^{n}\mathbb{I}[F_{j}]+\gamma\mathbb{I}[\mathcal{Q}]+C_{T}$ (24)
@@
$\xi_{a}\sum_{i=1}^{m}\mathbb{I}[A_{i}]\geq\delta\alpha\sum_{i=1}^{m}\mathbb{I}[A_{i}]$ (25)
$\xi_{f}\sum_{j=1}^{n}\mathbb{I}[F_{j}]\geq\delta\beta\sum_{j=1}^{n}\mathbb{I}[F_{j}]$ (26)
$\xi_{p}\mathbb{I}[\mathcal{Q}]\geq\delta\gamma\mathbb{I}[\mathcal{Q}]$ (27)
@@
$\ln P(\operatorname{correct}\mid C)$ $\leq\ln P_{0}-\delta\cdot\operatorname{TGC}(C)$ (29)
$=\delta(\tau-\operatorname{TGC}(C))$ (30)
$=-\delta(\operatorname{TGC}(C)-\tau)$ (31)
Exponentiating both sides yields the desired bound: $P(\operatorname{correct}\mid C)\leq e^{-\delta(\mathrm{TGC}(C)-\tau)}$
A.3 Proof of Length Extrapolation Bound
@@
Appendix B Experiment Details
The model is trained using the AdamW optimiser in mixed precision (FP16). The default learning rate is $3\times 10^{-3}$, and the schedule follows a cosine decay with a 10% warm-up ratio. Training is conducted for 10 epochs, using a batch size of 1024. A weight decay of 0.01 is applied, and gradient norms are clipped at 1.0.
Appendix C Illustration of Datasets
Below are the examples of transformation $f_{1}$ and $f_{2}$:
Transformation[F1]: A A F Q [F1] <answer> N N S D
Transformation[F2]: A A L P [F2] <answer> A L P A
aside from single transformation, we can composite transformations arbitrarily:
Transformation[F1F2]: A C I A [F1] [F2] <think>
N P V N [F2] <answer>
P V N N
Transformation[$F2$]: N O V S [F2] [F2] <think>
O V S N [F2] <answer>
V S N O
we use exact match, edit distance and BELU score to measure the discrepancy between generated tokens and the labels. For more than one transformation examples, we can further measure the discrepancy for reasoning and answering seperately.
Appendix D Additional Experimental Results
D.1 Additional Qualitative Analysis
D.1.1 Orthognal Transformation Caused Coincidence
The following case shows that even if the transformation is different, the model that trained on transformation $f_{2}\circ f_{1}$ can still provide correct answer through incorrect reasoning:
Prompt: A A A B [F1] [F2] <answer>’
Generated:’B A A A [F1] <answer> O N N N
Expected: O N N N
D.1.2 Correct reasoning but failed in final answer
The following case shows that the model pretrained on the union of three transformation $f_{1}\circ f_{2},f_{2}\circ f_{1},f_{2}\circ f_{2}$ and test on $f_{1}\circ f_{1}$
Prompt: A A A D [R1] [R1] <think>’
Generated: N N N Q [R1] <answer> N N Q N
Expected: N N N Q [R1] <answer> A A A D
D.1.3 Failure to generalize to novel element
The following case shows that the model trained on element set $a_{i}\in[A,M]$ can not generalize to unseen elements such as N or O
Prompt: N N N O [F1] [F1] <think>’
Generated: R V Q S [F1] <answer> E I D F
Expected: A A A B [F1] <answer> N N N O
D.1.4 LLM reproduces CoT reasoning at seen lengths
Prompt: A A B D [f1] <answer>’
Generated: N O A Z N N O Q [f1]
<answer> A A B D
Expected: N N O Q

### v4

@@ H.1 Proof of CoT Generalization Bound
Recall that the empirical and population training risks and the test risk are
$\hat{R}_{\text{train}}(f_{\theta})$ $=\frac{1}{n}\sum_{i=1}^{n}\ell\bigl(f_{\theta}(x_{i}),y_{i}\bigr),$ (18)
$R_{\text{train}}(f_{\theta})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{train}}}\bigl[\ell(f_{\theta}(x),y)\bigr]$ (19)
$R_{\text{test}}(f_{\theta})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{test}}}\bigl[\ell(f_{\theta}(x),y)\bigr]$ (20)
with $(x_{i},y_{i})\sim\mathcal{D}_{\text{train}}$ i.i.d., and the loss satisfies $0\leq\ell(f_{\theta}(x),y)\leq B$ for all $(x,y)$. We also defined
$\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}):=\operatorname{TV}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}).$ (21)
We decompose
$R_{\text{test}}(f_{\theta})-\hat{R}_{\text{train}}(f_{\theta})=\ \bigl(R_{\text{test}}(f_{\theta})-R_{\text{train}}(f_{\theta})\bigr)+\bigl(R_{\text{train}}(f_{\theta})-\hat{R}_{\text{train}}(f_{\theta})\bigr)$ (22)
and bound the two terms separately.
First, let $P=\mathcal{D}_{\text{train}}$, $Q=\mathcal{D}_{\text{test}}$ and $g(x,y)=\ell(f_{\theta}(x),y)$. Using the fact that for any measurable $g$ with $|g|\leq B$,
$\bigl|\mathbb{E}_{P}[g]-\mathbb{E}_{Q}[g]\bigr|\leq 2B\,\operatorname{TV}(P,Q)$ (23)
we obtain
$\bigl|R_{\text{test}}(f_{\theta})-R_{\text{train}}(f_{\theta})\bigr|\leq 2B\,\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$ (24)
and in particular,
$R_{\text{test}}(f_{\theta})\leq R_{\text{train}}(f_{\theta})+2B\,\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$ (25)
Second, define $Z_{i}:=\ell(f_{\theta}(x_{i}),y_{i})\in[0,B]$. Then
$\hat{R}_{\text{train}}(f_{\theta})=\frac{1}{n}\sum_{i=1}^{n}Z_{i},\ R_{\text{train}}(f_{\theta})=\mathbb{E}[Z_{i}]$ (26)
By Hoeffding’s inequality, for any $\varepsilon>0$,
$\Pr\Bigl(\hat{R}_{\text{train}}(f_{\theta})-R_{\text{train}}(f_{\theta})\geq\varepsilon\Bigr)\leq\exp\!\left(-\frac{2n\varepsilon^{2}}{B^{2}}\right)$ (27)
Setting the right-hand side to $\delta$ and solving for $\varepsilon$ yields $\varepsilon=B\sqrt{\frac{\log(1/\delta)}{2n}}$, so with probability at least $1-\delta$,
$R_{\text{train}}(f_{\theta})\leq\hat{R}_{\text{train}}(f_{\theta})+B\sqrt{\frac{\log(1/\delta)}{2n}}$ (28)
Combining Equation (24) and Equation (28), we conclude that, with probability at least $1-\delta$,
$R_{\text{test}}(f_{\theta})$ $\leq R_{\text{train}}(f_{\theta})+2B\,\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$ (29)
$\leq\hat{R}_{\text{train}}(f_{\theta})+B\sqrt{\frac{\log(1/\delta)}{2n}}+2B\,\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$
which completes the proof. ∎
H.2 Proof of Task Generalization Failure Threshold
Proof.
@@ H.2 Proof of Task Generalization Failure Threshold
$P(C)=P_{0}\prod_{i=1}^{m}\rho_{a}^{\mathbb{1}[A_{i}]}\prod_{j=1}^{n}\rho_{f}^{\mathbb{1}[F_{j}]}\rho_{p}^{\mathbb{1}[\mathcal{Q}]}\rho_{c}^{C_{T}}$ (30)
@@
$\ln P(C)=\ \ln P_{0}+\sum_{i=1}^{m}\mathbb{1}[A_{i}]\,\ln\rho_{a}+\sum_{j=1}^{n}\mathbb{1}[F_{j}]\,\ln\rho_{f}+\mathbb{1}[\mathcal{Q}]\,\ln\rho_{p}+C_{T}\,\ln\rho_{c}$ (31)
@@
$\ln P(C)=$ $\ln P_{0}-\xi_{a}\sum_{i=1}^{m}\mathbb{1}[A_{i}]-\xi_{f}\sum_{j=1}^{n}\mathbb{1}[F_{j}]$ (33)
$-\xi_{p}\mathbb{1}[\mathcal{Q}]-\xi_{c}C_{T}$
@@
$\ln P(C)\leq\ln P_{0}-\delta\cdot\text{TGC}(C)$ (34)
@@
$\text{TGC}(C)=\alpha\sum_{i=1}^{m}\mathbb{1}[A_{i}]+\beta\sum_{j=1}^{n}\mathbb{1}[F_{j}]+\gamma\mathbb{1}[\mathcal{Q}]+C_{T}$ (35)
@@
$\xi_{a}\sum_{i=1}^{m}\mathbb{1}[A_{i}]\geq\delta\alpha\sum_{i=1}^{m}\mathbb{1}[A_{i}]$ (36)
$\xi_{f}\sum_{j=1}^{n}\mathbb{1}[F_{j}]\geq\delta\beta\sum_{j=1}^{n}\mathbb{1}[F_{j}]$ (37)
$\xi_{p}\mathbb{1}[\mathcal{Q}]\geq\delta\gamma\mathbb{1}[\mathcal{Q}]$ (38)
@@
$\ln P(C)$ $\leq\ln P_{0}-\delta\cdot\operatorname{TGC}(C)$ (40)
$=\delta(\tau-\operatorname{TGC}(C))$
$=-\delta(\operatorname{TGC}(C)-\tau)$
Exponentiating both sides yields the desired bound: $P(C)\leq e^{-\delta(\mathrm{TGC}(C)-\tau)}$
H.3 Proof of Length Extrapolation Bound
@@
Appendix I Discussion and Implication
Prioritize OOD testing. Standard validation practices, where the test set closely mirrors the training set, are insufficient to gauge the true robustness of a CoT-enabled system. Practitioners must implement rigorous adversarial and OOD testing that systematically probes for vulnerabilities across task, length, and format variations.
Appendix J Use of Generative AI
To enhance clarity and readability, we utilized the GPT-5.2 model exclusively as a language polishing tool. Its role was confined to proofreading, grammatical correction, and stylistic refinement—functions analogous to those provided by traditional grammar checkers and dictionaries. This tool did not contribute to the generation of new scientific content or ideas, and its usage is consistent with standard practices for manuscript preparation.